A cone has a diameter of 16 units and a height of 11 units. How many cubic units is the volume of the cone?


1
Expert's answer
2022-04-15T14:16:42-0400

The formula of volume for cone is :

V="\frac{SH}{3}" = "\frac{64•11\pi}{3}" =235"\pi" cubic units~737 cubic units